# Ultrafast Heating Induced Suppression of d-band Dominance in the Electronic Excitation Spectrum of Cuprum

### v1

@@ 1 Introduction
In this way, it was shown that the laser-induced heating of electrons leads to the lattice instability and melting (disordering) of silicon due to the weakening of the inter-ionic bonds 20, 21. Another example is the laser-induced solid–solid phase transition that has been observed in bismuth 22. In contrast to semiconductors, metals can remain stable under laser heating of the electrons, and, strikingly, can even manifest a more rigid lattice structure. For example, Descamps et al. 17 have recently reported the observation of a stable gold crystal lattice where the electrons have been heated by the FEL to a few electronvolts. In this experiment, a signature of phonon hardening has been observed, whereby the bonds stiffen between atoms. This effect was earlier predicted by ab initio Kohn-Sham density functional theory (KS-DFT) calculations 23 of the hot electrons within the cold ionic lattice using the local density approximation (LDA) for the exchange–correlation (XC) functional. A second example for the successful utilization of KS-DFT for predicting the properties of solids with laser-excited electrons is the calculation of the XRTS spectrum of an isochorically heated aluminium foil by Mo et al. 24 based on linear-response time-dependent DFT (LR-TDDFT) using an adiabatic LDA (ALDA) XC kernel. The same combination of LR-TDDFT with ALDA was shown to accurately describe the XRTS spectrum of aluminum at ambient conditions, where the plasmon was measured with ultrahigh resolution at the European XFEL 25. Therefore, LR-TDDFT can be expected to yield accurate results for the electronic dynamic structure factor of isochorically heated metals across temperature regimes.
@@
Very recently, Moldabekov et al. 26 have used this approach to study the effect of electronic heating on the order of a few electronvolts on the expected XRTS spectrum 26; this has revealed an interesting red shift of the plasmon energy by 0.10.10.1\,0.1eV for aluminium and by 111\,1eV for silicon as a consequence of thermal excitations. In the case of aluminium, the effect is small and only manifests at small wavenumbers $q\lesssim 0.1{\rm\AA^{-1}}$, making it very challenging to measure. For silicon, on the other hand, the plasmon shift of $1~{}{\rm eV}$ at temperatures $T\simeq 2{\rm eV}$ is well within experimental measurement capabilities 27, 25. However, the possible instability of the lattice due to the weakening of the inter-ionic bonds 20, 21 can be a serious obstacle in practice. Therefore, it is important to ask if such a heating-induced red shift prominently manifests itself in other metals that are stable under FEL radiation. Going back to aluminium, an additional thermally induced feature is the formation of a double plasmon peak as the region of Landau damping is shifted to lower wavenumbers upon increasing the electronic temperature 26. This effect is similar to the formation of the double plasmon in the DSF of ground-state aluminum near the pair continuum 28, 29, 30. These results show that thermal excitations in X-ray-driven solids can generate a variety of new features in the XRTS spectrum at a finite momentum transfer.
@@
In the present work, we carry out extensive new LR-TDDFT calculations to explore the XRTS spectrum of isochorically heated copper. In contrast to simple metals, the effect of d-states dominates over plasmon-type excitations in the DSF of electrons in transition metals 31, 32. In gold and copper, excitations originating in the d-band lead to the formation of a prominent double peak structure at $\omega>\omega_{p}$ and a substantial broadening of the plasmon feature at $\omega=\omega_{p}$. Interestingly, the presence of the d-state excitations leads to a plasmon dispersion that is nearly independent of the wavenumber for both materials 31, 32. Here, we investigate in detail the interplay of these effects with thermal excitations on the DSF of copper at different temperatures, wavenumbers, and crystallographic directions. Indeed, thermal effects on the DSF are profound: we find an emerging collective plasmon excitation that becomes dominant over the $d$-band feature for $T\gtrsim 4\,$eV and which starts to follow the familiar Bohm-Gross relation in this regime. In addition, we find a pronounced blue-shift of the plasmon with increasing $T$, which is in stark contrast to other isochorically heated metals such as Al 26. Finally, we discuss the possibility to use XRTS experiments with isochorically heated copper as a rigorous testbed for the theoretical modeling of warm dense matter (WDM) 33, 34, 35—an extreme state that occurs in astrophysical objects 36, 3, 37, 38, and which plays an important role e.g. for inertial confinement fusion 4, 39, 6 applications.
@@ 2.1 Theoretical Framework
The intensity that is measured in an XRTS experiment is given by a convolution of the electronic dynamic structure factor $S(\mathbf{q},\omega)$ with the combined source-and-instrument function $R(\omega_{s})$ 12,
$I(\mathbf{q},\omega_{s})=S(\mathbf{q},\omega)\circledast R(\omega_{s})\ ,$ (1)
where the latter accounts both for the finite width of the probing X-ray source, and for all effects of the detector 40. The momentum transfer $\mathbf{q}$ is determined from the scattering angle and $\omega=\omega_{0}-\omega_{s}$ denotes the energy loss of the scattered photon where $\omega_{0}$ and $\omega_{s}$ are the initial beam energy and the actual energy of the photon captured by the detector. The state-of-the-art is given by the European XFEL in Germany, where XRTS measurements with the capability of resolving electronic features with a resolution of up to $\Delta\omega\sim 0.1\,{\rm eV}$ have been recently demonstrated 25.
To study the effect of thermal electronic excitations on the XRTS spectrum of X-ray-driven copper, we use the LR-TDDFT method with an adiabatic XC kernel. Indeed, LR-TDDFT constitutes the most common method to study the DSF of solids, and there is a vast body of dedicated literature, see e.g.Ref. 41 and references therein; here, we restrict ourselves to a concise overview of the main ideas.
As a first step, we consider the well-known fluctuation–dissipation theorem that connects the macroscopic dielectric function $\epsilon_{M}(\mathbf{q},\omega)$ with $S(\mathbf{q},\omega)$ 42, 43,
@@
where $n$ denotes the electronic number density, and $e$ the elementary charge. The term “macroscopic” indicates that $\epsilon_{M}(\mathbf{q},\omega)$ describes the volume averaged response to an external perturbation 44, 45, 46. It is computed by taking the diagonal part of the inverse microscopic dielectric matrix $\epsilon_{M}^{-1}(\mathbf{q},\omega)=\left[\epsilon^{-1}(\mathbf{k},\omega)\right]_{\mathbf{G}\mathbf{G}}$, where $\mathbf{q}=\mathbf{G}+\mathbf{k}$ (with $\mathbf{k}$ being in the first Brillouin zone) and $\mathbf{G}$ is a reciprocal lattice vector 47, 48. The latter is defined by the microscopic density response function  41,
@@
The LR-TDDFT method allows one to compute $\chi_{\scriptscriptstyle\mathbf{G},\mathbf{G}^{\prime}}(\mathbf{k},\omega)$ in different approximations. The lowest rank corresponds to the so-called independent particle approximation (IPA). In the IPA, the Kohn-Sham (KS) orbitals and eigenenergies are used to calculate the density response function $\chi^{~{}0}_{\scriptscriptstyle\mathbf{G},\mathbf{G}^{\prime}}(\mathbf{k},\omega)$ according to the ideal electron gas model 49. Since the KS eigenenergies from the self-consistent ground state (equilibrium state) calculations are employed, $\chi^{~{}0}_{\scriptscriptstyle\mathbf{G},\mathbf{G}^{\prime}}(\mathbf{k},\omega)$ already has information about excitations between different orbitals. However, being computed using a formula for the ideal Fermi gas model, $\chi^{~{}0}_{\scriptscriptstyle\mathbf{G},\mathbf{G}^{\prime}}(\mathbf{k},\omega)$ omits various correlation effects, such as screening due to the Hartree mean field, microscopic density inhomogeneities due to the field of the ions etc. The inclusion of correlation effects leads to a Dyson’s type equation for the density response function $\chi_{\scriptscriptstyle\mathbf{G},\mathbf{G}^{\prime}}(\mathbf{k},\omega)$ 50, 48:
@@
where $v_{\scriptscriptstyle\mathbf{G}1}(\mathbf{k})={4\pi}/{|\mathbf{k}+\mathbf{G}_{1}|^{2}}$ is the Coulomb potential in reciprocal-space, and $K^{\rm xc}_{\scriptscriptstyle\mathbf{G}_{1},\mathbf{G}_{2}}(\mathbf{k},\omega)$ is the XC kernel capturing electronic correlations; it is defined as the functional derivative of the XC potential in KS-DFT 51.
In our calculations, we have used a static (adiabatic) XC kernel $K^{\rm xc}_{\scriptscriptstyle\mathbf{G}_{1},\mathbf{G}_{2}}(\mathbf{k},\omega=0)$ within the ALDA 41. The ALDA is known to provide a fairly accurate description of the macroscopic dielectric function $\epsilon_{M}(\mathbf{q},\omega)$ of metals and semiconductors at finite wavenumbers 25, 52, 53, 30, 54. The relevant thermal signatures explored in this work are characterized by a difference of $\delta\omega\gtrsim 1~{}{\rm eV}$ from the ground state features. If needed, a further fine-tuning can be achieved either by employing more advanced static XC kernels beyond ALDA 55, 56 or by using an explicitly dynamic approximation (e.g., see Ref. 50 and references therein) in future works.
We note that, in the ground state, the DSF is usually studied indirectly by measuring the electronic energy loss spectrum (EELS), see, e.g., Refs. 57, 58, 59. In principle, the XRTS spectrum and the EELS are directly related since
@@
From Eq. (2), one can see that $S(\mathbf{q},\omega)\sim q^{2}~{}{\rm EELS}(\mathbf{q},\omega)$. This means that EELS is advantageous for measurements at small wavenumbers, whereas XRTS might be more suitable at large wavenumbers. However, EELS measurements are problematic for experiments with matter under extreme conditions due to its requirements for thin targets as well as long measurement times 59, which are not realistic for the transient states that are of interest in the current work.
@@ 2.2 Calculation parameters
We used the GPAW code 60, 61, 62, 63, 64, 65, which is a real-space implementation of the projector augmented-wave (PAW) approach 66. We used the ground-state LDA XC-functional by Perdew and Wang 67. The simulations have been carried out for a face-centered cubic (fcc) lattice with the lattice parameter $3.61$ set according to the experimental value 68. For the calculation of the KS states, we used the energy cutoff $1000~{}{\rm eV}$, the PAW dataset of copper provided by GPAW, and the primitive cell combined with the $k$-point grid $40\times 40\times 40$. We note that in the employed LR-TDDFT formalism, the momentum transfer must be the difference between two $k$-points times $2\pi/a$. To study the possible impact of inhomogeneity with respect to crystallographic directions on the DSF, calculations were performed along the [100], [111], and [011] directions. We considered electronic temperatures in the range $0.025~{}{\rm eV}\leq T\leq 12~{}\rm{eV}$ and the number of KS bands was set to $N_{b}=100$. The smearing of the occupation numbers was computed according to the Fermi-Dirac distribution. On the stage of the calculation of the density response matrix, the local field effect cutoff was set to $150~{}{\rm eV}$. In all calculations, we used $\eta=0.1~{}{\rm eV}$ for the Lorentzian smearing parameter in $\chi_{\scriptscriptstyle\mathbf{G},\mathbf{G}^{\prime}}(\mathbf{k},\omega)$ 63.
3 Simulations Results
@@ 3.1 DSF $S(\mathbf{q},\omega)$ in the X-ray driven copper
We start our investigation by considering the electronic ground state (here represented by the results for $T=0.025\,$eV) in the limit of small wave numbers. In the left panel of Fig. 1, we show the EELS spectrum at $q=0.0754$-1 for different temperatures in the range from 0.025 eV up to 8 eV (with the temperature increasing in the subplots from bottom to top). We focus on four main features of the EELS spectrum (which equivalently appear in the DSF, cf. Fig. 4) denoted by capital letters A–D. A thorough investigation of the EELS properties of copper in the ground state has been presented by Alkauskas et al 31, where it was shown that features A and B are plasmon type collective oscillations, whereas C and D are a consequence of excitations between the d-band and the unoccupied states above the Fermi level. More specifically, peak B can be described as a collective plasmon oscillations of the valence electrons; this has been shown by Campillo et al 32 by freezing the d-band into the core and leaving only the $4s^{1}$ state that forms the valence electron. Our LR-TDDFT results are in good agreement with the structure of electronic excitations reported in Refs. 31, 32, and we reproduce the positions of all four peaks A–D.
@@
Following the analysis by Campillo et al 32 who investigated the DSF of copper at ambient conditions, signatures B and C can be understood by considering the density of states (DOS), which is shown in the right panel of Fig. 1 as the solid lines. The $d$-band dominates the accumulation of the states below the Fermi energy. This is shown by the projected density of states on $d$-states, which is depicted by the shaded area. We find that, with increasing temperatures, the d-states are shifted to lower energies. Since the features C and D in the EELS/DSF emerge as the result of the transitions from d-states to high-lying bands above the Fermi level, the observed shift of the d-states results in a blue-shift by approximately the same amount.
@@
An additional interesting research topic is due to the lattice structure, which is known to lead to an anisotropy of the DSF with respect to the crystallographic direction at certain wavenumbers 69, 31. To quantify this anisotropy effect on X-ray driven copper, we show the DSF in the [011] direction at different temperatures and wavenumbers in Fig. 4. We find that the DSF in the [011] direction closely resembles the DSF in the [100] direction for $q\lesssim 0.9$-1, and differences in the shape and peak positions emerge for $q\gtrsim 1.2$-1. For completeness, we note that the DSF in [111] direction is equivalent to the [100] direction (see Appendix). In summary, we conclude that the electronic DSF of copper is nearly isotropic at $T\geq 4~{}{\rm eV}$ for $q\lesssim 0.9$-1.
@@
Figure 5: The demonstration of the role of the factor $f(\omega)=\left(1-e^{-\hbar\omega/k_{B}T}\right)^{-1}$ on the enhancement of the DSF features at small energies. Solid lines are the results for the DSF along the [100] direction and dashed lines are the same data divided by $f(\omega)$.

### v2

@@ 1 Introduction
In this way, it was shown that the laser-induced heating of electrons leads to the lattice instability and melting (disordering) of silicon due to the weakening of the inter-ionic bonds 20, 21. In contrast to semiconductors, metals can remain stable under laser heating of the electrons, and, strikingly, can even manifest a more rigid lattice structure. For example, Descamps et al. 17 have recently reported the observation of a stable gold crystal lattice where the electrons have been heated by the FEL to a few electronvolts. In this experiment, a signature of phonon hardening has been observed, whereby the bonds stiffen between atoms. This effect was earlier predicted by ab initio Kohn-Sham density functional theory (KS-DFT) calculations 22 of the hot electrons within the cold ionic lattice using the local density approximation (LDA) for the exchange–correlation (XC) functional. A second example for the successful utilization of KS-DFT for predicting the properties of solids with laser-excited electrons is the calculation of the XRTS spectrum of an isochorically heated aluminium foil by Mo et al. 23 based on linear-response time-dependent DFT (LR-TDDFT) using an adiabatic LDA (ALDA) XC kernel. The same combination of LR-TDDFT with ALDA was shown to accurately describe the XRTS spectrum of aluminum at ambient conditions, where the plasmon was measured with ultrahigh resolution at the European XFEL 24. Therefore, LR-TDDFT can be expected to yield accurate results for the electronic dynamic structure factor of isochorically heated metals across temperature regimes.
@@
Very recently, Moldabekov et al. 25 have used this approach to study the effect of electronic heating on the order of a few electronvolts on the expected XRTS spectrum 25; this has revealed an interesting red shift of the plasmon energy by 0.10.10.1\,0.1eV for aluminium and by 111\,1eV for silicon as a consequence of thermal excitations. In the case of aluminium, the effect is small and only manifests at small wavenumbers $q\lesssim 0.1{\rm\AA^{-1}}$, making it very challenging to measure. For silicon, on the other hand, the plasmon shift of $1~{}{\rm eV}$ at temperatures $T\simeq 2{\rm eV}$ is well within experimental measurement capabilities 26, 24. However, the possible instability of the lattice due to the weakening of the inter-ionic bonds 20, 21 can be a serious obstacle in practice. Therefore, it is important to ask if such a heating-induced red shift prominently manifests itself in other metals that are stable under FEL radiation. Going back to aluminium, an additional thermally induced feature is the formation of a double plasmon peak as the region of Landau damping is shifted to lower wavenumbers upon increasing the electronic temperature 25. This effect is similar to the formation of the double plasmon in the DSF of ground-state aluminum near the pair continuum 27, 28, 29. These results show that thermal excitations in X-ray-driven solids can generate a variety of new features in the XRTS spectrum at a finite momentum transfer.
@@
In the present work, we carry out extensive new LR-TDDFT calculations to explore the XRTS spectrum of isochorically heated copper. In contrast to simple metals, the effect of d-states dominates over plasmon-type excitations in the DSF of electrons in transition metals 30, 31. In gold and copper, excitations originating in the d-band lead to the formation of a prominent double peak structure at $\omega>\omega_{p}$ and a substantial broadening of the plasmon feature at $\omega=\omega_{p}$. Interestingly, the presence of the d-state excitations leads to a plasmon dispersion that is nearly independent of the wavenumber for both materials 30, 31. Here, we investigate in detail the interplay of these effects with thermal excitations on the DSF of copper at different temperatures, wavenumbers, and crystallographic directions. Indeed, thermal effects on the DSF are profound: we find an emerging collective plasmon excitation that becomes dominant over the $d$-band feature for $T\gtrsim 4\,$eV and which starts to follow the familiar Bohm-Gross relation in this regime. In addition, we find a pronounced blue-shift of the plasmon with increasing $T$, which is in stark contrast to other isochorically heated metals such as Al 25. Finally, we discuss the possibility to use XRTS experiments with isochorically heated copper as a rigorous testbed for the theoretical modeling of warm dense matter (WDM) 32, 33, 34—an extreme state that occurs in astrophysical objects 35, 3, 36, 37, and which plays an important role e.g. for inertial confinement fusion 4, 38, 6 applications.
@@ 2.1 Theoretical Framework
The intensity that is measured in an XRTS experiment is given by a convolution of the combined source-and-instrument function $R(\omega_{s})$ 12 and the electronic dynamic structure factor $S(\mathbf{q},\omega)=S(\mathbf{q},-\Delta\omega)$ (with $\Delta\omega$ being the energy loss of the scattered photon),
$I(\mathbf{q},\Delta\omega_{s})=S(\mathbf{q},\Delta\omega)\circledast R(\omega_{s})\ ,$ (1)
where the latter accounts both for the finite width of the probing X-ray source, and for all effects of the detector 39. The momentum transfer $\mathbf{q}$ is determined from the scattering angle. The state-of-the-art is given by the European XFEL in Germany, where XRTS measurements with the capability of resolving electronic features with a resolution of up to $\delta\omega\sim 0.1\,{\rm eV}$ have been recently demonstrated 24.
To study the effect of thermal electronic excitations on the XRTS spectrum of X-ray-driven copper, we use the LR-TDDFT method with an adiabatic XC kernel. Indeed, LR-TDDFT constitutes the most common method to study the DSF of solids, and there is a vast body of dedicated literature, see e.g.Ref. 40 and references therein; here, we restrict ourselves to a concise overview of the main ideas.
As a first step, we consider the well-known fluctuation–dissipation theorem that connects the macroscopic dielectric function $\epsilon_{M}(\mathbf{q},\omega)$ with $S(\mathbf{q},\omega)$ 41, 42,
@@
where $n$ denotes the electronic number density, and $e$ the elementary charge. The term “macroscopic” indicates that $\epsilon_{M}(\mathbf{q},\omega)$ describes the volume averaged response to an external perturbation 43, 44, 45. It is computed by taking the diagonal part of the inverse microscopic dielectric matrix $\epsilon_{M}^{-1}(\mathbf{q},\omega)=\left[\epsilon^{-1}(\mathbf{k},\omega)\right]_{\mathbf{G}\mathbf{G}}$, where $\mathbf{q}=\mathbf{G}+\mathbf{k}$ (with $\mathbf{k}$ being in the first Brillouin zone) and $\mathbf{G}$ is a reciprocal lattice vector 46, 47. The latter is defined by the microscopic density response function  40,
@@
The LR-TDDFT method allows one to compute $\chi_{\scriptscriptstyle\mathbf{G},\mathbf{G}^{\prime}}(\mathbf{k},\omega)$ in different approximations. The lowest rank corresponds to the so-called independent particle approximation (IPA). In the IPA, the Kohn-Sham (KS) orbitals and eigenenergies are used to calculate the density response function $\chi^{~{}0}_{\scriptscriptstyle\mathbf{G},\mathbf{G}^{\prime}}(\mathbf{k},\omega)$ according to the ideal electron gas model 48. Since the KS eigenenergies from the self-consistent ground state (equilibrium state) calculations are employed, $\chi^{~{}0}_{\scriptscriptstyle\mathbf{G},\mathbf{G}^{\prime}}(\mathbf{k},\omega)$ already has information about excitations between different orbitals. However, being computed using a formula for the ideal Fermi gas model, $\chi^{~{}0}_{\scriptscriptstyle\mathbf{G},\mathbf{G}^{\prime}}(\mathbf{k},\omega)$ omits various correlation effects, such as screening due to the Hartree mean field, microscopic density inhomogeneities due to the field of the ions etc. The inclusion of correlation effects leads to a Dyson’s type equation for the density response function $\chi_{\scriptscriptstyle\mathbf{G},\mathbf{G}^{\prime}}(\mathbf{k},\omega)$ 49, 47:
@@
where $v_{\scriptscriptstyle\mathbf{G}1}(\mathbf{k})={4\pi}/{|\mathbf{k}+\mathbf{G}_{1}|^{2}}$ is the Coulomb potential in reciprocal-space, and $K^{\rm xc}_{\scriptscriptstyle\mathbf{G}_{1},\mathbf{G}_{2}}(\mathbf{k},\omega)$ is the XC kernel capturing electronic correlations; it is defined as the functional derivative of the XC potential in KS-DFT 50.
The LR-TDDFT method provides the DSF of the electrons in the thermal equilibrium. An alternative approach that can perform the simulation of electronic dynamics with a distribution different from the Fermi-Dirac distribution is real-time TDDFT (RT-TDDFT), where electronic wave functions are propagated according to time-dependent KS equations. This method was used by Silaeva et al. 51 to study ultrafast electron dynamics thermalization in metals driven by a 7-fs laser pulse. Silaeva et al.51 showed that valence electrons reach a thermalized state within the time of the laser pulse. The RT-TDDFT method can also be used to compute the DSF. For example, Baczewski et al. 52 used RT-TDDFT to compute the DSF of warm dense beryllium in thermal equilibrium. We note that if the same XC functional were used in both, RT-TDDFT and LR-TDDFT are formally equivalent for linear response properties in thermal equilibrium  40.
In our calculations, we have used a static (adiabatic) XC kernel $K^{\rm xc}_{\scriptscriptstyle\mathbf{G}_{1},\mathbf{G}_{2}}(\mathbf{k},\omega=0)$ within the ALDA 40. The ALDA is known to provide a fairly accurate description of the macroscopic dielectric function $\epsilon_{M}(\mathbf{q},\omega)$ of metals and semiconductors at finite wavenumbers 24, 53, 54, 29, 55. The relevant thermal signatures explored in this work are characterized by a difference of $\delta\omega\gtrsim 1~{}{\rm eV}$ from the ground state features. If needed, a further fine-tuning can be achieved either by employing more advanced static XC kernels beyond ALDA 56, 57 or by using an explicitly dynamic approximation (e.g., see Ref. 49 and references therein) in future works.
We note that, in the ground state, the DSF is usually studied indirectly by measuring the electronic energy loss spectrum (EELS), see, e.g., Refs. 58, 59, 60. In principle, the XRTS spectrum and the EELS are directly related since
@@
From Eq. (2), one can see that $S(\mathbf{q},\omega)\sim q^{2}~{}{\rm EELS}(\mathbf{q},\omega)$. This means that EELS is advantageous for measurements at small wavenumbers, whereas XRTS might be more suitable at large wavenumbers. However, EELS measurements are problematic for experiments with matter under extreme conditions due to its requirements for thin targets as well as long measurement times 60, which are not realistic for the transient states that are of interest in the current work.
In consistency with measurements at FEL facilities, we consider the electronic response on subpicosecond time scales and treat the ions as being frozen in their crystal lattice positions surrounded by heated electrons. This is justified since the electron-lattice equilibration time is order of picoseconds 61, 62, 63, 64, 65, 66. Furthermore, this approximation is corroborated by the predictions of increased melting temperature with electron heating in copper and other d-band metals 67, 64, and by the recent observation of phonon hardening in gold 17.
@@ 2.2 Calculation parameters
We used the GPAW code 68, 69, 70, 71, 72, 73, which is a real-space implementation of the projector augmented-wave (PAW) approach 74. We used the ground-state LDA XC-functional by Perdew and Wang 75. The simulations have been carried out for a face-centered cubic (fcc) lattice with the lattice parameter $3.61$ set according to the experimental value 76. For the calculation of the KS states, we used the energy cutoff $1000~{}{\rm eV}$, the PAW dataset of copper provided by GPAW (with $1s-3p$ orbitals treated as frozen core electrons), and the primitive cell combined with the $k$-point grid $40\times 40\times 40$. We note that in the employed LR-TDDFT formalism, the momentum transfer must be the difference between two $k$-points times $2\pi/a$. To study the possible impact of inhomogeneity with respect to crystallographic directions on the DSF, calculations were performed along the [100], [111], and [011] directions. We considered electronic temperatures in the range $0.025~{}{\rm eV}\leq T\leq 12~{}\rm{eV}$ and the number of KS bands was set to $N_{b}=100$. The smearing of the occupation numbers was computed according to the Fermi-Dirac distribution. On the stage of the calculation of the density response matrix, the local field effect cutoff was set to $150~{}{\rm eV}$. In all calculations, we used $\eta=0.1~{}{\rm eV}$ for the Lorentzian smearing parameter in $\chi_{\scriptscriptstyle\mathbf{G},\mathbf{G}^{\prime}}(\mathbf{k},\omega)$ 71. For the density of state (DOS) calculations of copper, we used the same parameters as for the DSF. The DOS was plotted by setting the Gaussian width parameter to $0.2~{}{\rm eV}$.
3 Simulations Results and Discussion
@@ 3.1 DSF $S(\mathbf{q},\omega)$ in the X-ray driven copper
We start our investigation by considering the electronic ground state (here represented by the results for $T=0.025\,$eV) in the limit of small wave numbers. In the left panel of Fig. 1, we show the EELS spectrum at $q=0.0754$-1 for different temperatures in the range from 0.025 eV up to 8 eV (with the temperature increasing in the subplots from bottom to top). We focus on four main features of the EELS spectrum (which equivalently appear in the DSF, cf. Fig. 4) denoted by capital letters A–D. A thorough investigation of the EELS properties of copper in the ground state has been presented by Alkauskas et al 30, where it was shown that features A and B are plasmon type collective oscillations, whereas C and D are a consequence of excitations between the d-band and the unoccupied states above the Fermi level. More specifically, peak B can be described as a collective plasmon oscillations of the valence electrons; this has been shown by Campillo et al 31 by freezing the d-band into the core and leaving only the $4s^{1}$ state that forms the valence electron. Our LR-TDDFT results are in good agreement with the structure of electronic excitations reported in Refs. 30, 31, and we reproduce the positions of all four peaks A–D.
@@
Following the analysis by Campillo et al 31 who investigated the DSF of copper at ambient conditions, signatures B and C can be understood by considering the density of states (DOS), which is shown in the right panel of Fig. 1 as the solid lines. The $d$-band dominates the accumulation of the states below the Fermi energy. This is shown by the projected density of states on $d$-states, which is depicted by the shaded area. We find that, with increasing temperatures, the d-states are shifted to lower energies. Since the features C and D in the EELS/DSF emerge as the result of the transitions from d-states to high-lying bands above the Fermi level, the observed shift of the d-states results in a blue-shift by approximately the same amount. An accurate determination of the difference in the shifts of the C and D peaks is difficult due to the strong broadening of the C peak. Nevertheless, one can observe that the blue-shift of C and D peaks have close values at $T<4~{}{\rm eV}$. We estimate that the D peak of the EELS shifts by about $1~{}{\rm eV}$ more towards larger frequencies at $T=8~{}{\rm eV}$ and $T=6~{}{\rm eV}$ compared to the C peak.
@@
An additional interesting research topic is due to the lattice structure, which is known to lead to an anisotropy of the DSF with respect to the crystallographic direction at certain wavenumbers 77, 30. To quantify this anisotropy effect on X-ray driven copper, we show the DSF in the [011] direction at different temperatures and wavenumbers in Fig. 4. We find that the DSF in the [011] direction closely resembles the DSF in the [100] direction for $q\lesssim 0.9$-1, and differences in the shape and peak positions emerge for $q\gtrsim 1.2$-1. For completeness, we note that the DSF in [111] direction is equivalent to the [100] direction (see Appendix). In summary, we conclude that the electronic DSF of copper is nearly isotropic at $T\geq 2~{}{\rm eV}$ for $q\lesssim 0.9$-1.
@@
Figure 5: The demonstration of the role of the factor $f(\omega)=\left(1-e^{-\hbar\omega/k_{B}T}\right)^{-1}$ on the enhancement of the DSF features at small energies. Solid lines are the results for the DSF along the [100] direction and dashed lines are the same data divided by $f(\omega)$. The wavenumber values are given in the units of Å-1.